Freshman Seminar 21n: Elliptic Curve Projects

William Stein

April 22, 2003

Today you will hand in your first draft. Your assignment for the week is to work on a more polished draft of your project and send me an email, with one paragraph about the presentations today (including your own). A more polished draft of your project is due Tuesday, April 29.

Schedule

April 29 Fourth presentation (a more polished draft is due)
May 6 Fifth presentation (get comments on your draft back)
May 13 Give a polished 40 minute presentation and final version

(Note: I will be gone May 5–6 (at IDA), but Grigor will lead class those two days.)